Find four consecutive even integers such that twice the sum of the first and fourth is 124
Answer by tutorcecilia(2152)

Find four consecutive even integers such that twice the sum of the first and fourth is 124
1st=x
2nd=x+2
3rd=x+4
4th=x+6
.
so,
2[(x)+(x+6)]=124
2(2x+6)=124
4x+12=124
4x=124-12
4x=112
x=28
so,
1st =28
2nd=30
3rd=32
4th=34
(2)(28+34)=124
124=124